# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 7 Twisted projections of tubes satisfying the Convex Wolff Axioms
Let $I\subset\mathbb{R}$ be a compact interval and let $M\subset C^{2}(I)$ be a family of cinematic functions (see [16, Definition 1.6]). Then for all $\varepsilon>0$, there exists $\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $G\subset M$ be a $\delta$-separated set that satisfies the Frostman non-concentration condition
@@
Let $X_{0}=\underset{T\in\mathbb{T}}{\bigcup}\pi_{f}(Y(T))$ and $0<\eta\ll\varepsilon$ be a small parameter to be chosen later. For each dyadic number $1<\lambda<\delta^{-4}$, let $X_{\lambda}$ be the set of $x\in X_{0}$ such that $\sum_{p\in\pi_{f}^{-1}(x)}\#\mathbb{T}(p)\sim\lambda$, where the sum is over a $\delta$-separated set of $p\in\pi_{f}^{-1}(x)$ and #$\#\mathbb{T}(p)$ was defined in (6.1). By pigeonholing, there exists $\lambda$ such that $Y_{\lambda}(T):=Y(T)\cap\pi_{f}^{-1}(X_{\lambda})$ is a $|\log\delta|^{-2}\delta^{\eta}$-dense shading.
@@
After possibly refinement and rotation, each tube $T\in\mathbb{T}$ is $\delta$-neighborhood of a line of the form $l_{a,b,c,d}$ intersecting the unit cube. Let $A(\mathbb{T})$ denote set of parameters of these lines, i.e.
@@
Since $\mathbb{T}$ satisfies the Tube Wolff axioms, if $\rho\in[\delta,1]$ then every $\rho$-tube $T_{\rho}$ contains at most $\delta^{-\eta}|T_{\rho}|(\#\mathbb{T})$ tubes from $\mathbb{T}$, and hence the point set $A(\mathbb{T})\subset\mathbb{R}^{4}$ of parameters is a $(\delta,2,\delta^{-\eta})$-Frostman set.
Let $A_{c}=\{(a,b,d):(a,b,c,d)\in A(\mathbb{T})\}$. By Fubini, there exists a $c\in[0,1]$ such that $\mathcal{E}_{\delta}(A_{c}):=\delta^{-s}\geq\delta^{-1+\eta}\geq\delta^{-1+\varepsilon^{2}}$. Apply Lemma 7.8 to refine $A_{c}$ to be an $\eta$-uniform subset.
So far we only know the $\delta$-covering number of $A_{c}$ but not the spacing condition. To apply Theorem 7.12, we need to know the spacing of $A_{c}$. By Lemma 7.9, there exists $\rho^{\prime}\in(\delta^{1-\varepsilon^{2}},1]$ such that for any $\rho$-cube $Q\subset N_{\rho^{\prime}}(A_{c})$, $A_{Q}:=S_{Q}(A_{c}\cap Q)$ is a $(\rho,1-\varepsilon^{2},\rho^{-12\varepsilon^{2}})$-set with $\rho=\delta/\rho^{\prime}$.
By Lemma 7.6, since $A_{Q}$ is a $(\rho,1-\varepsilon^{2},\rho^{-12\varepsilon^{2}})$-set, there exists a $(\rho,1-\varepsilon^{2},100)$-Katz-Tao set $A_{Q}^{\prime}\subset A_{Q}$ with $\mathcal{E}_{\rho}(A_{Q}^{\prime})\gtrsim\rho^{14\varepsilon^{2}-1}.$ Then $S_{Q}^{-1}(A_{Q}^{\prime})$ is a $(\delta,1-\varepsilon^{2},100)$-Katz-Tao set contained in $Q$, a cube of side-length $\rho^{\prime}=\delta/\rho$. By Lemma 7.10, there exist $\sim\rho^{\prime-1}$ translations $A_{i}$ such that the set $A_{c}^{\prime}:=\sqcup_{i}A_{i}(S_{Q}^{-1}(A_{Q}^{\prime}))$ is a $(\rho,1-\varepsilon^{2},100)$-Katz-Tao set with $\mathcal{E}_{\delta}(A_{c}^{\prime})\gtrsim\rho^{14\varepsilon^{2}}\delta^{-1}.$ For each $A_{i}$ and $(a,b,d)\in S_{Q}^{-1}(A_{Q}^{\prime})$, define the shading on the tube $T^{\prime}=T_{a^{\prime},b^{\prime},c,d^{\prime}}$, where $(a^{\prime},b^{\prime},d^{\prime})=A_{i}(a,b,d)$, as
@@
where the union $\bigcup_{A_{i}}$ is taken over all $(a,b,d)\in A_{i}(S_{Q}^{-1}(A_{Q}^{\prime}))$ and the corresponding $T=T_{a,b,c,d}$. To see that (7.4) holds, let $A_{j}\circ A_{i}^{-1}(x)=x+(a_{0},b_{0},d_{0})$. Then $a+ct+f(t)(b+dt)\in\Omega_{i,t}:=\cup_{A_{i}}\pi_{f,t}(Y^{\prime}(T))$ if and only if
$a+a_{0}+ct+f(t)(b+b_{0}+(d+d_{0})t)\in\Omega_{j,t}:=\cup_{A_{j}}\pi_{f,t}(Y^{\prime}(T)).$
@@
Note that the LHS of (7.6) is contained in the $10\delta/\rho$-neighborhood of a curve $\pi_{f}(l_{a,b,c,d})$, $(a,b,d)\in Q$, whose volume is comparable to the RHS of (7.6), up to a factor of $\rho^{-45\varepsilon^{2}}$. We conclude that the $\eta$-uniform set $X=\bigcup_{T\in\mathbb{T}}\pi_{f}(Y^{\prime}(T))$ satisfies
@@
Let $0<\eta\ll\varepsilon$ to be determined. Let $\mathbb{T}^{\prime}\subset\mathbb{T}$ be the set of tubes with $|Y(T)|\geq\delta^{2\eta}|T|$, then #$\#\mathbb{T}^{\prime}\geq\delta^{\eta}\#\mathbb{T}.$ Apply Lemma 7.8 to find an $\eta$-uniform subset $\mathbb{T}^{\prime\prime}\subset\mathbb{T}^{\prime}$ with #$\#\mathbb{T}^{\prime\prime}\geq\delta^{\eta}\mathbb{T}^{\prime}$. Then $(\mathbb{T}^{\prime\prime},Y)_{\delta}$ becomes a set of $\delta$-tubes satisfying the Tube Wolff Axioms with error $\delta^{-3\eta}$ and $Y$ is a $\delta^{2\eta}$-dense shading. To ease the notation, still let $\mathbb{T}$ denote $\mathbb{T}^{\prime\prime}$.
@@
$\Big{|}\bigcup_{T\in\mathbb{T}}\pi_{f}(Y(T))\Big{|}\geq\Big{|}\bigcup_{T\in\mathbb{T}}\pi_{f}(Y_{1}(T))\Big{|}\geq(\delta/\rho_{1})^{\varepsilon}\Big{|}N_{\rho_{1}}\big{(}\bigcup_{T\in\mathbb{T}}\pi_{f}(Y_{1}(T))\big{)}\Big{|}.$
Consider $\mathbb{T}_{\rho_{1}}$ as the set of distinct $\rho_{1}$-tubes containing at least one tube in $\mathbb{T}$. Since $\mathbb{T}$ is $\eta$-uniform, $\mathbb{T}_{\rho_{1}}$ satisfies the Tube Wolff Axioms with error $\delta^{-3\eta}$. Define $Y_{\rho_{1}}$ be a shading on $\tilde{T}\in\mathbb{T}_{\rho_{1}}$ as $Y_{\rho_{1}}=N_{\rho_{1}}(\cup_{T\in\mathbb{T}[\tilde{T}]}Y_{1}(T))$. Therefore, $Y_{\rho_{1}}$ is a $\delta^{2\eta}$-dense shading. Apply Lemma 7.11 on $(\mathbb{T}_{\rho_{1}},Y_{\rho_{1}})_{\rho_{1}}$ and find $\rho_{2}\in(\rho_{1}^{1-\varepsilon^{2}},1)$ and corresponding $Y_{2}\subset Y_{\rho_{1}}$ such that
@@
Since for any $\tilde{T}\in\mathbb{T}_{\rho_{1}}$, $Y_{2}\subset Y_{\rho_{1}}\subset N_{\rho_{1}}(\cup_{T\in\mathbb{T}[\tilde{T}]}\pi_{f}(Y_{1}(T))$, we have
@@
Suppose to the contrary that there exists a Kakeya set $K$ with essentially equal Hausdorff and packing dimension $\alpha<3$. Without loss of generality, we can assume $K\subset B(0,1)$. Let $\varepsilon=(3-\alpha)/4$; in what follows, all implicit constant may depend on $\varepsilon$. Let $\eta>0$ and $\delta_{0}>0$ be the quantities from Theorem 1.5 for this choice of $\varepsilon$. Let $\eta_{1}=(3-\varepsilon)\eta/6$. Since $K$ has essentially equal Hausdorff and packing dimension, we can find a measure $\mu$ supported on $K$ so that
@@
We have $|Y(T_{e})|\gtrsim\tau|T|$ for each tube $T_{e}$; we will select $\delta\in(0,\delta_{0}]$ sufficiently small so that $|Y(T_{e})|\geq\delta^{\eta}|T|$. Let $\mathbb{T}=\{T_{e}\colon e\in\Omega_{\delta}\}$. Since the tubes in $\mathbb{T}$ point in $\delta$-separated directions, the tubes are essentially distinct and satisfy the Convex Wolff Axioms with error $O(\delta^{-2}/(\#\mathbb{T}))=O(|\Omega|^{-1})$; we will select $\delta\in(0,\delta_{0}]$ sufficiently small so that $\mathbb{T}$ satisfies the convex Wolff Axioms with error at most $\delta^{-\eta}$. Thus we can apply Theorem 1.5 to conclude there exist scales $\rho,r\in[\delta,1]$ with $\rho\leq\delta^{\eta}r$, and a ball $B$ of radius $r$ so that (1.1) holds.
@@
If $\delta\in(0,\delta_{0}]$ is selected sufficiently small, then this is impossible. We conclude that every Kakeya set in $\mathbb{R}^{3}$ with essentially equal Hausdorff and packing dimension must have dimension 3. ∎
@@
After dyadic pigeonholing, we can select $\rho,r\in[\delta,1]$ with $\rho\leq\delta^{\eta}r$, and a set of indices $I$ with the following properties.
  • For each index $i\in I$ we have $\rho_{i}\in[\rho,2\rho)$ and $r_{i}\in[r,2r)$
  • $\sum_{T\in\mathbb{T}}\Big{|}Y(T)\cap\bigcup_{i\in I}B_{i}\Big{|}\gtrsim|\log\delta|^{-2}\sum_{T\in\mathbb{T}}|Y(T)|.$
Next, we can greedily select a set of indices $I^{\prime}\subset I$ so that the above properties continue to hold (with the implicit constant in the second item worsened slightly), so that the sets $\{B_{i}\colon i\in I^{\prime}\}$ are pairwise disjoint.
We claim that the values of $\rho$ and $r$ described above, plus the shading $Y^{\prime}(T)=Y(T)\cap\bigcup_{i\in I^{\prime}}B_{i}$ satisfies the conclusions of Corollary 1.7. Indeed, we can compute

### v2

@@ 7 Twisted projections of tubes satisfying the Convex Wolff Axioms
Let $I\subset\mathbb{R}$ be a compact interval and let $M\subset C^{2}(I)$ be a family of cinematic functions (see [16, Definition 1.6]). Then for all $\varepsilon>0$, there exists $\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $G\subset M$ be a $\delta$-separated set that satisfies the following Frostman non-concentration condition:
@@
Let $X_{0}=\bigcup_{\mathbb{T}}\pi_{f}(Y(T))$, and let $0<\eta\ll\varepsilon$ be a small parameter to be chosen later. For each dyadic number $1<\lambda<\delta^{-4}$, let $X_{\lambda}$ be the set of $x\in X_{0}$ such that $\sum_{p\in\pi_{f}^{-1}(x)}\#\mathbb{T}(p)\sim\lambda$, where the sum is over a $\delta$-separated set of $p\in\pi_{f}^{-1}(x)$ and #$\#\mathbb{T}(p)$ was defined in (6.1). By pigeonholing, there exists $\lambda$ such that $Y_{\lambda}(T):=Y(T)\cap\pi_{f}^{-1}(X_{\lambda})$ is a $|\log\delta|^{-2}\delta^{\eta}$-dense shading.
@@
After possibly refinement and rotation, each tube $T\in\mathbb{T}$ is $\delta$-neighborhood of a line of the form $l_{a,b,c,d}$ intersecting the unit cube. Let $A(\mathbb{T})$ denote the set of tuples $(a,b,c,d)$ corresponding to these lines, i.e.
@@
Since $\mathbb{T}$ satisfies the Tube Wolff axioms, by Remark 7.4 we have that $A(\mathbb{T})\subset\mathbb{R}^{4}$ is a $(\delta,2,\delta^{-\eta})$-Frostman set.
Let $A_{c}=\{(a,b,d):(a,b,c,d)\in A(\mathbb{T})\}$. By Fubini, there exists a $c\in[0,1]$ such that $\mathcal{E}_{\delta}(A_{c}):=\delta^{-s}\geq\delta^{-1+\eta}\geq\delta^{-1+\varepsilon^{2}}$. Apply Lemma 7.9 to replace $A_{c}$ with an $\eta$-uniform subset, which we will continue to denote by $A_{c}$.
At this point we have a bound on the $\delta$-covering number of $A_{c}$, but we do not know if this set satisfies our desired non-concentration condition. To apply Theorem 7.13, we need to know that $A_{c}$ satisfies (7.3). By Lemma 7.10, there exists $\rho^{\prime}\in(\delta^{1-\varepsilon^{2}},1]$ such that for any $\rho$-cube $Q\subset N_{\rho^{\prime}}(A_{c})$, $A_{Q}:=S_{Q}(A_{c}\cap Q)$ is a $(\rho,1-\varepsilon^{2},\rho^{-12\varepsilon^{2}})$-set with $\rho=\delta/\rho^{\prime}$.
By Lemma 7.6, since $A_{Q}$ is a $(\rho,1-\varepsilon^{2},\rho^{-12\varepsilon^{2}})$-set, there exists a $(\rho,1-\varepsilon^{2},100)$-Katz-Tao set $A_{Q}^{\prime}\subset A_{Q}$ with $\mathcal{E}_{\rho}(A_{Q}^{\prime})\gtrsim\rho^{14\varepsilon^{2}-1}.$ Then $S_{Q}^{-1}(A_{Q}^{\prime})$ is a $(\delta,1-\varepsilon^{2},100)$-Katz-Tao set contained in $Q$, a cube of side-length $\rho^{\prime}=\delta/\rho$. By Lemma 7.11, there exist $\sim\rho^{\prime-1}$ translations $A_{i}$ such that the set $A_{c}^{\prime}:=\bigsqcup_{i}A_{i}(S_{Q}^{-1}(A_{Q}^{\prime}))$ is a $(\rho,1-\varepsilon^{2},100)$-Katz-Tao set with $\mathcal{E}_{\delta}(A_{c}^{\prime})\gtrsim\rho^{14\varepsilon^{2}}\delta^{-1}.$ For each $A_{i}$ and $(a,b,d)\in S_{Q}^{-1}(A_{Q}^{\prime})$, define the shading on the tube $T^{\prime}=T_{a^{\prime},b^{\prime},c,d^{\prime}}$, where $(a^{\prime},b^{\prime},d^{\prime})=A_{i}(a,b,d)$, as
@@
where the union $\bigcup_{A_{i}}$ is taken over all $(a,b,d)\in A_{i}(S_{Q}^{-1}(A_{Q}^{\prime}))$ and the corresponding $T=T_{a,b,c,d}$. To see that (7.4) holds, let $A_{j}\circ A_{i}^{-1}(x)=x+(a_{0},b_{0},d_{0})$. Then $a+ct+f(t)(b+dt)\in\Omega_{i,t}:=\bigcup_{A_{i}}\pi_{f,t}(Y^{\prime}(T))$ if and only if
$a+a_{0}+ct+f(t)(b+b_{0}+(d+d_{0})t)\in\Omega_{j,t}:=\bigcup_{A_{j}}\pi_{f,t}(Y^{\prime}(T)).$
@@
Note that the LHS of (7.6) is contained in the $10\delta/\rho$-neighborhood of a curve $\pi_{f}(l_{a,b,c,d})$, $(a,b,d)\in Q$, whose volume is comparable to the RHS of (7.6), up to a factor of $\rho^{-45\varepsilon^{2}}$. We conclude that the $\eta$-uniform set $X=\bigcup_{\mathbb{T}}\pi_{f}(Y^{\prime}(T))$ satisfies
@@
Let $0<\eta\ll\varepsilon$ be a small number to be determined below. Let $\mathbb{T}^{\prime}\subset\mathbb{T}$ be the set of tubes with $|Y(T)|\geq\delta^{2\eta}|T|$; then #$\#\mathbb{T}^{\prime}\geq\delta^{\eta}(\#\mathbb{T}).$ Apply Lemma 7.9 to find an $\eta$-uniform subset $\mathbb{T}^{\prime\prime}\subset\mathbb{T}^{\prime}$ with #$\#\mathbb{T}^{\prime\prime}\geq\delta^{\eta}(\#\mathbb{T}^{\prime})$. Then $(\mathbb{T}^{\prime\prime},Y)_{\delta}$ becomes a set of $\delta$-tubes satisfying the Tube Wolff Axioms with error $\delta^{-3\eta}$, and $Y$ is a $\delta^{2\eta}$-dense shading. To ease notation, we will rename $\mathbb{T}^{\prime\prime}$ to be $\mathbb{T}$.
@@
$\Big{|}\bigcup_{T\in\mathbb{T}}\pi_{f}(Y(T))\Big{|}\geq\Big{|}\bigcup_{T\in\mathbb{T}}\pi_{f}(Y_{1}(T))\Big{|}\geq(\delta/\rho_{1})^{\varepsilon}\Big{|}N_{\rho_{1}}\big{(}\bigcup_{T\in\mathbb{T}}\pi_{f}(Y_{1}(T))\big{)}\Big{|}.$ (7.7)
Consider $\mathbb{T}_{\rho_{1}}$ as the set of distinct $\rho_{1}$-tubes containing at least one tube in $\mathbb{T}$. Since $\mathbb{T}$ is $\eta$-uniform, $\mathbb{T}_{\rho_{1}}$ satisfies the Tube Wolff Axioms with error $\delta^{-3\eta}$. Define $Y_{\rho_{1}}$ to be the shading on $\tilde{T}\in\mathbb{T}_{\rho_{1}}$ given by $Y_{\rho_{1}}=N_{\rho_{1}}\big{(}\bigcup_{\mathbb{T}[\tilde{T}]}Y_{1}(T)\big{)}$. Therefore, $Y_{\rho_{1}}$ is a $\delta^{2\eta}$-dense shading. Apply Lemma 7.12 to $(\mathbb{T}_{\rho_{1}},Y_{\rho_{1}})_{\rho_{1}}$ and find $\rho_{2}\in(\rho_{1}^{1-\varepsilon^{2}},1)$ and corresponding $Y_{2}\subset Y_{\rho_{1}}$ such that
@@
Since for any $\tilde{T}\in\mathbb{T}_{\rho_{1}}$, we have
$Y_{2}(\tilde{T})\subset Y_{\rho_{1}}(\tilde{T})=N_{\rho_{1}}\Big{(}\bigcup_{T\in\mathbb{T}[\tilde{T}]}Y_{1}(T)\Big{)},$
we have
@@
Suppose to the contrary that there exists a Kakeya set $K$ with stably equal Hausdorff and packing dimension $\alpha<3$. Without loss of generality, we can assume $K\subset B(0,1)$. Let $\varepsilon=(3-\alpha)/4$; in what follows, all implicit constant may depend on $\varepsilon$. Let $\eta>0$ and $\delta_{0}>0$ be the quantities from Theorem 1.5 for this choice of $\varepsilon$. Let $\eta_{1}=(3-\varepsilon)\eta/6$. Since $K$ has stably equal Hausdorff and packing dimension, we can find a measure $\mu$ supported on $K$ so that
@@
We have $|Y(T_{e})|\gtrsim\tau|T|$ for each tube $T_{e}$; we will select $\delta\in(0,\delta_{0}]$ sufficiently small so that $|Y(T_{e})|\geq\delta^{\eta}|T|$. Let $\mathbb{T}=\{T_{e}\colon e\in\Omega_{\delta}\}$. Since the tubes in $\mathbb{T}$ point in $\delta$-separated directions, the tubes are essentially distinct and satisfy the Convex Wolff Axioms with error $O(\delta^{-2}/(\#\mathbb{T}))=O(|\Omega|^{-1})$; we will select $\delta\in(0,\delta_{0}]$ sufficiently small so that $\mathbb{T}$ satisfies the convex Wolff Axioms with error at most $\delta^{-\eta}$. Thus we can apply Theorem 1.5 to conclude that there exist scales $\rho,r\in[\delta,1]$ with $\rho\leq\delta^{\eta}r$, and a ball $B$ of radius $r$ so that (1.1) holds.
@@
If $\delta\in(0,\delta_{0}]$ is selected sufficiently small, then this is impossible. We conclude that every Kakeya set in $\mathbb{R}^{3}$ with stably equal Hausdorff and packing dimension must have dimension 3. ∎
@@
After dyadic pigeonholing, we can select $\rho,r\in[\delta,1]$ with $\rho\leq\delta^{\eta}r$, and a set of indices $I$ with the following properties:
  • For each index $i\in I$, we have $\rho_{i}\in[\rho,2\rho)$ and $r_{i}\in[r,2r)$.
  • $\sum_{T\in\mathbb{T}}\Big{|}Y(T)\cap\bigcup_{i\in I}B_{i}\Big{|}\gtrsim|\log\delta|^{-2}\sum_{\mathbb{T}}|Y(T)|.$
Next, we can greedily select a set of indices $I^{\prime}\subset I$ so that the above properties continue to hold (with the implicit constant in the second item worsened slightly), and the sets $\{B_{i}\colon i\in I^{\prime}\}$ are pairwise disjoint.
We claim that the values of $\rho$ and $r$ described above, plus the shading $Y^{\prime}(T)=Y(T)\cap\bigcup_{i\in I^{\prime}}B_{i}$ satisfy the conclusions of Corollary 1.7. Indeed, we can compute